You are the owner of a parasailing company that is expanding operations to a new beachfront location, and you need to prepare a three-year analysis for the bank that may loan you the funds to purchase your boat and parasailing equipment. Because of your well-established reputation, you already have received requests for “flights” to be scheduled as soon as you open the new location. Therefore, you expect to break-even the first year but must calculate the number of flights needed. You also need to determine the new break-even point in Year 2 if the location allows referrals, which you believe will average about 2% of the sales price overall. Finally, you need to determine the volume needed to have $10,000 in profit in Year 3. The following information is available: Sales price per flight $175 Estimated loan payment per month $350 Fuel costs per flight $100 Full-time scheduler salary $2,500 per month Boat crew per flight $30 $500 per month dock fee and use of a small office on the pier Requirements: Calculate the Year 1 break-even quantity, contribution margin, and contribution margin ratio. Explain how the values were determined. Calculate the Year 2 break-even quantity, break-even sales, and contribution margin ratio. Explain how the values were determined. Determine the number of flights (units) needed to retain a profit of $10,000 in Year 3, assuming the company does allow for referrals. Recommend if the bank should issue the loan. Superior papers will: Perform all calculations correctly. Articulate the approach to solving the problem. Explain the relationship of the costs to the concept of contribution margin. Discuss any limitations of the data, including what may be missing. Conclude on whether the bank should issue the loan
Answer #1

Here we have to find

1) Contribution Margin and

3) Contribution margin ratio

3 ) Break Even Quantity

1) CONTRIBUTION MARGIN Formula

   SALE PRICE PER UNIT - VARIABLE COST PER UNIT ,

SALE PRICE $ 175 - VARIABLE COST ( i.e fuel Cost ) 100 = $ 75

THUS , CONTRIBUTION MARGIN IS $ 75 OF YEAR 1

NOW CONTRIBUTION MARGIN FOR YEAR TWO IS

Here it is written that in year 2 Sale price and variable cost will be increase by 2 % so

NEW SALE PRICE IS 175 + 175*2/100 = $ 178.5

NEW VARIABLE COST IS 100 + 100*2/100= $ 102

YEAR 2 CONTRIBUTION WILL BE

SALE PRICE PER UNIT - VARIABLE COST PER UNIT  

=178.5-100

= $ 76.5

2 ) CONTRIBUTION MARGIN RATIO FORMULA

TOTAL SALE PER UNIT - VARIABLE COST PER UNIT

------------------------------------------------------------------------------ * 100

TOTAL SALE PER UNIT

YEAR 1 CONTRIBUTION MARGIN RATIO =

(SALE PER UNIT-VARIABLE COST PER UNIT / SALE PER UNIT )*100

(175-100/175)*100

= 42.85 %

YEAR 2 CONTRIBUTION RATION

SALE PER UNIT-VARIABLE COST PER UNIT / SALE PER UNIT

178.5-102/178.5*100

= 42.86 %

3 ) BREAK EVEN QUANTITY

FIX COST / CONTRIBUTION MARGIN PER UNIT

FIX COST WILL BE SAME FOR YEAR 1 AND YEAR 2

FIX COST = LOAN PAYMENT PER MONTH $ 350

+ FULL TIME SALARY $ 2500

+ BOAT CREW PER FLIGHT $ 30

+ DOCK FEE $ 500

--------------------

TOTAL $ 3380

SO 1 YEAR BREAK EVEN UNITS =  

FIX COST / CONTRIBUTION PER UNITS

= 3380 /75 ( $ 75 ( i.e. calculated in point no 1)

= 45.06 UNITS

SO 2 YEAR BREAK EVEN UNITS =  

FIX COST / CONTRIBUTION PER UNITS

= 3380 /76.5 ( $ 76.5 i.e. calculated in point no 1)

= 44.18 UNITS

NOW BREAK EVEN SALES'S CALCULATIONS

= BREAK EVEN UNITS * SALE PRICE

FOR YEAR 1 BREAK EVEN SALE

=45.06*175

= $ 7885.50

FOR YEAR 2 BREAK EVEN SALE

= 44.18*178.5

= $ 7885.13